OpenStudy (anonymous):

What is F(X)=X^2+3X-2? If the given domain is F(-1)

OpenStudy (anonymous):

The domain of inverse will be the range of the orginal functions

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

Thankks;)

OpenStudy (hoblos):

F(-1)=(-1)^2 +3(-1) -2 =1-3-2 =-4